# Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness

### v1

@@ Appendix A Methodology
This section provides comprehensive algorithmic details and theoretical proofs for the SCER framework presented in the main paper. We first present detailed step-by-step procedures for the SCER implementation, including covariance computation, embedding regularization, and loss optimization that were summarized in the main text. Subsequently, we provide formal proofs for the theoretical foundations underlying our embedding regularization approach, including the decomposition of the embedding space into core and spurious components and the optimality conditions for robustness.
@@ A.1 Overall process of SCER
The Spurious Loss $\mathcal{L}_{\text{spur}}$ penalizes alignment with spurious directions, scaled by the magnitude of spurious variation in the data. The Core Loss $\mathcal{L}_{\text{core}}$ encourages alignment with core directions, with the negative sign ensuring that stronger core alignment reduces the overall loss. Our loss formulation penalizes alignment with spurious directions and encourages alignment with core directions. By introducing control parameters $\lambda_{\text{spur}}$ and $\lambda_{\text{core}}$, we derive our embedding loss function:
@@
0:  Data $(x,y)$, parameters $(w,\beta)$, hyperparams $(\eta,\lambda_{\text{spur}},\lambda_{\text{core}})$
@@ A.2 Theoretical Analysis of SCER
We now present a theoretical analysis that formally decomposes the worst-group error. Following prior work (yao2022improving), we adopt a Gaussian mixture model to analyze how worst-group error arises from label-domain interactions.
@@ A.2.1 Binary Setting
Consider a classification problem where each data point $x\in\mathcal{X}$ is associated with a label $\mathcal{Y}=\{y_{-1},y_{+1}\}$ and a domain $\mathcal{D}=\{d_{R},d_{G}\}$. Each pair $(y,d)\in\mathcal{Y}\times\mathcal{D}$ defines a subpopulation. We assume that the data follows a group-conditional Gaussian distribution $x\mid(y,d)\sim\mathcal{N}(\mu^{(y,d)},\Sigma)$ where $\Sigma$ is positive definite. We further assume that $\Delta_{\text{core}}=\mu^{(y_{+1},d)}-\mu^{(y_{-1},d)}$ is constant across $d$, $\Delta_{\text{spur}}=\mu^{(y,d_{R})}-\mu^{(y,d_{G})}$ is constant across $y$, $\Sigma^{(d_{R})}=\Sigma^{(d_{G})}=\Sigma$, $\mathbb{P}(y=\pm 1)=\mathbb{P}(d\in\{d_{R},d_{G}\})=\frac{1}{2}$, and $\mathbb{P}_{y,d}\approx 0$ for some pairs $(y,d)$ under extreme spurious correlations.
@@
$\log\frac{p(x\mid y_{-1})}{p(x\mid y_{+1})}=(\mu^{(y_{-1})}-\mu^{(y_{+1})})^{\top}\Sigma^{-1}x-\tfrac{1}{2}(\mu^{(y_{-1})}+\mu^{(y_{+1})})^{\top}\Sigma^{-1}(\mu^{(y_{-1})}-\mu^{(y_{+1})}),$
which is linear in $x$ with a normal vector $\Sigma^{-1}(\mu^{(y_{-1})}-\mu^{(y_{+1})})$. Since logistic regression is Fisher-consistent under correct model specification, the cross-entropy minimizer $\beta^{*}$ aligns with the Bayes direction:
$\beta^{*}\propto\Sigma^{-1}(\mu^{(y_{-1})}-\mu^{(y_{+1})}).$
In the extreme shift setting, the effective class means correspond to $\mu^{(y_{-1},d_{R})}$ and $\mu^{(y_{+1},d_{G})}$. Thus
$\beta^{*}\propto\Sigma^{-1}\big(\mu^{(y_{-1},d_{R})}-\mu^{(y_{+1},d_{G})}\big).$
@@
$\mu^{(y_{-1},d_{R})}-\mu^{(y_{+1},d_{G})}=(\mu^{(y_{-1},d_{R})}-\mu^{(y_{+1},d_{R})})+(\mu^{(y_{+1},d_{R})}-\mu^{(y_{+1},d_{G})})=\Delta_{\mathrm{core}}+\Delta_{\mathrm{spur}}=\tilde{\Delta}$,
@@
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\operatorname{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-\operatorname{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right),$
@@
Using $\beta_{0}=-\beta^{\top}\mathbb{E}[x]$ under extreme spurious correlations, we
$E(y_{-1},d_{R})$ $=\Phi\!\left(-\tfrac{1}{2}\tfrac{\tilde{\Delta}^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$
$E(y_{+1},d_{R})$ $=\Phi\!\left(\tfrac{-\Delta_{\mathrm{core}}^{\top}\beta+\tfrac{1}{2}\tilde{\Delta}^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$
$E(y_{-1},d_{G})$ $=\Phi\!\left(\tfrac{(\tfrac{1}{2}\tilde{\Delta}-\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$
$E(y_{+1},d_{G})$ $=\Phi\!\left(\tfrac{\Delta_{\mathrm{core}}^{\top}\beta-\tfrac{1}{2}\tilde{\Delta}^{\top}\beta}{\|\beta\|_{\Sigma}}\right).$
@@
$E(y_{-1},d_{R})=\Phi(-z),\qquad E(y_{+1},d_{G})=\Phi(z).$
Since $\Phi(-z)=1-\Phi(z)$, one of these exceeds $0.5$ while the other is below $0.5$. The seen subgroups $(y_{-1},d_{R}),(y_{+1},d_{G})$ have arguments in [z,z][-z,z], hence their errors lie in $\Phi(-z),\Phi(z)]$.
$E_{\mathrm{wge}}=\max\{\Phi(-z),\Phi(z)\}.$
Finally, writing $\tilde{\Delta}=\Delta_{\mathrm{spur}}+\Delta_{\mathrm{core}}$ gives
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\frac{\beta^{\top}\Delta_{\mathrm{spur}}}{\|\beta\|_{\Sigma}}-\frac{\beta^{\top}\Delta_{\mathrm{core}}}{\|\beta\|_{\Sigma}}\right).$
With Proposition 1 and the identity
@@
The sign $\pm$ appearing in Theorem 1 originates from the fact that the ERM solution Under cross-entropy loss is determined only up to an overall sign. That is, both $\Sigma^{-1}\tilde{\Delta}$ and $\Sigma^{-1}\tilde{\Delta}$ define equivalent solutions, since the classifier depends only on the direction of $\beta^{*}$. Consequently, the $\pm$ in the expression of the worst-group error does not alter the actual decision boundary or error behavior, but we retain it explicitly to make the sign ambiguity transparent.
@@
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\mathrm{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-\mathrm{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right),$ (4)
@@
Equation (4) shows that the worst-group error is governed by two competing components:
@@
The first term increases the error through alignment with spurious directions, while the second decreases the error through alignment with core directions. Thus, minimizing $E_{\mathrm{wge}}$ requires simultaneously suppressing the spurious term and enhancing the core term.
@@
In Appendix 1, we derived Eq. (4) under the strict ERM setting with Gaussian assumptions, where the optimal $\beta^{*}$ has the closed form $\beta^{*}\propto\Sigma^{-1}\tilde{\Delta}$. In practice, however, $\beta$ corresponds to the weight parameters of a deep neural classifier trained via stochastic optimization. While $\beta$ may not coincide with the ERM minimizer, the decomposition in Eq. (4) remains structurally valid: worst-group error is still determined by the trade-off between spurious and core alignment. This observation justifies applying the embedding regularization loss in Eq. (5) to deep models.
A.2.2 Multiclass/Multidomain Setting
We now generalize this assumption to multiclass, multidomain settings to investigate whether analogous results are preserved.
Consider a classification problem where each data point $x\in\mathcal{X}$ is associated with a label $\mathcal{Y}=\{y_{1},\dots,y_{m}\}$ and a domain $\mathcal{D}=\{d_{1},\dots,d_{k}\}$. Each pair $(y,d)\in\mathcal{Y}\times\mathcal{D}$ defines a subpopulation. We assume that the data follows a group-conditional Gaussian distribution $x\mid(y,d)\sim\mathcal{N}(\mu^{(y,d)},\Sigma)$ where $\Sigma$ is positive definite. We further assume that $\Delta_{\text{core}}^{(i,j)}=\mu^{(y_{i},d)}-\mu^{(y_{j},d)}$ is constant across $d$, $\Delta_{\text{spur}}^{(i,j)}=\mu^{(y,d_{i})}-\mu^{(y,d_{j})}$ is constant across $y$, $\Sigma^{(d_{1})}=\Sigma^{(d_{2})}=\cdots=\Sigma^{(d_{k})}=\Sigma$, $\mathbb{P}(y=y_{i})=\frac{1}{m},\;\mathbb{P}(d=d_{j})=\frac{1}{k}$, and $\mathbb{P}_{y,d}\approx 0$ for some pairs $(y,d)$ under extreme spurious correlations, meaning that most samples from a given label class come from a specific domain.
Proposition 2 (Multiclass/Multidomain ERM under Cross-Entropy Loss).
Under the assumptions, the ERM solution with multinomial logistic regression satisfies, for all $i\neq j$,
$\beta_{i}^{*}-\beta_{j}^{*}\ \propto\ \Sigma^{-1}\tilde{\Delta}^{(i,j)},\qquad\tilde{\Delta}^{(i,j)}=\Delta_{\mathrm{core}}^{(i,j)}+\Delta_{\mathrm{spur}}^{(p,q)},$
where $d_{p}$ and $d_{q}$ denote the predominant domains for classes $y_{i}$ and $y_{j}$, respectively, under the extreme spurious-correlation assumption.
Proof.
The ERM problem using cross-entropy loss for multinomial logistic regression is
$\min_{\{(\beta_{i},\beta_{0i})\}_{i=1}^{m}}\;\mathbb{E}\!\left[-\sum_{i=1}^{m}\mathbf{1}\{y=i\}\log\frac{\exp(\beta_{i}^{\top}x+\beta_{0i})}{\sum_{r=1}^{m}\exp(\beta_{r}^{\top}x+\beta_{0r})}\right].$
Assume the group-conditional Gaussian model with shared covariance,
$x_{\text{emb}}\mid(y=i,d=j)\sim\mathcal{N}\!\big(\mu^{(y_{i},d_{j})},\,\Sigma\big),$
where $\mu^{(y_{i},d_{j})}=\mathbb{E}_{x\sim\mathbb{P}_{y_{i},d_{j}}}[f_{w}(x)]$ denotes the mean embedding of subpopulation $(y_{i},d_{j})$. Let $\pi_{i}=\mathbb{P}(y=i)$ denote the class prior. The Bayes discriminant for class $i$ is
$\delta_{i}(x_{\text{emb}})=x_{\text{emb}}^{\top}\Sigma^{-1}\bar{\mu}_{i}-\tfrac{1}{2}\bar{\mu}_{i}^{\top}\Sigma^{-1}\bar{\mu}_{i}+\log\pi_{i},$
where $\delta_{i}(\cdot)$ denotes the LDA-type discriminant and $\bar{\mu}_{i}$ the effective class mean. Then, for any $i\neq j$,
$\log\frac{p(y=i\mid x_{\text{emb}})}{p(y=j\mid x_{\text{emb}})}=(\bar{\mu}_{i}-\bar{\mu}_{j})^{\top}\Sigma^{-1}x_{\text{emb}}-\tfrac{1}{2}\!\left(\bar{\mu}_{i}^{\top}\Sigma^{-1}\bar{\mu}_{i}-\bar{\mu}_{j}^{\top}\Sigma^{-1}\bar{\mu}_{j}\right)+\log\pi_{i}-\log\pi_{j},$
which is linear in $x_{\text{emb}}$ with normal vector $\Sigma^{-1}(\bar{\mu}_{i}-\bar{\mu}_{j})$. By Fisher consistency of multinomial logistic regression under the correctly specified linear logit model,
$\beta_{i}^{*}-\beta_{j}^{*}\ \propto\ \Sigma^{-1}(\bar{\mu}_{i}-\bar{\mu}_{j}).$
Under the extreme spurious-correlation assumption, each class $y_{i}$ predominantly appears in a single domain $d_{p}$, hence
$\bar{\mu}_{i}\approx\mu^{(y_{i},d_{p})}.$
Based on this, we decompose
$\bar{\mu}_{i}-\bar{\mu}_{j}\approx\big(\mu^{(y_{i},d_{p})}-\mu^{(y_{j},d_{p})}\big)+\big(\mu^{(y_{j},d_{p})}-\mu^{(y_{j},d_{q})}\big)=\Delta_{\mathrm{core}}^{(i,j)}+\Delta_{\mathrm{spur}}^{(p,q)}.$
Substituting into the previous relation gives
$\beta_{i}^{*}-\beta_{j}^{*}\ \propto\ \Sigma^{-1}\!\left(\Delta_{\mathrm{core}}^{(i,j)}+\Delta_{\mathrm{spur}}^{(p,q)}\right).$
Theorem 2 (Multiclass/Multidomain Worst-Group error).
Under these assumptions, for the softmax classifier be parameterized by $\{(\beta_{i},\beta_{0i})\}_{i=1}^{m}$ and define the pairwise standardized margin on subgroup $(y_{i},d_{p})$ against class $i\neq j$ as
$m_{i\to j}(\beta):=\frac{(\beta_{i}-\beta_{j})^{\top}\mu^{(y_{i},d_{p})}+(\beta_{0i}-\beta_{0j})}{\|\beta_{i}-\beta_{j}\|_{\Sigma}}.$
Then the subgroup error is bounded as
$\max_{i\neq j}\,\Phi\!\big(-m_{i\to j}(\beta)\big)\;\leq\;\mathbb{P}\!\big(f(x_{\text{emb}})\neq i\,\big|\,y=y_{i},\,d=d_{p}\big)\;\leq\;\sum_{i\neq j}\,\Phi\!\big(-m_{i\to j}(\beta)\big),$
and the worst-group error over observed subgroups satisfies
$\max_{i}\max_{i\neq j}\,\Phi\!\big(-m_{i\to j}(\beta)\big)\;\leq\;E_{\mathrm{wge}}(\beta)\;\leq\;\max_{i}\sum_{i\neq j}\,\Phi\!\big(-m_{i\to j}(\beta)\big).$
Proof.
Fix $(y=y_{i},d=d_{p})$. Since $x_{\text{emb}}\sim\mathcal{N}(\mu^{(y_{i},d_{p})},\Sigma)$, the pairwise score difference
$T_{i\to j}(x_{\text{emb}}):=(\beta_{i}-\beta_{j})^{\top}x_{\text{emb}}+(\beta_{0i}-\beta_{0j})$
is Gaussian with mean $(\beta_{i}-\beta_{j})^{\top}\mu^{(y_{i},d_{p})}+(\beta_{0i}-\beta_{0j})$ and variance $\|\beta_{i}-\beta_{j}\|_{\Sigma}^{2}$. Hence
$\mathbb{P}(\text{misclassify $i$ as $j$}\mid y=y_{i},d=d_{p})=\Phi(-m_{i\to j}(\beta)).$
The lower bound follows from $\mathbb{P}(\cup_{i\neq j}A_{j})\geq\max_{j}\mathbb{P}(A_{j})$ and the upper bound from the union bound $\mathbb{P}(\cup_{i\neq j}A_{j})\leq\sum_{j}\mathbb{P}(A_{j})$. ∎
From Theoretical Analysis to Embedding Loss (Multiclass/Multidomain)
Under Proposition 2, the population ERM solution satisfies $\beta_{i}^{*}-\beta_{j}^{*}\propto\Sigma^{-1}\tilde{\Delta}_{i,j}$ with $\tilde{\Delta}_{i,j}:=\Delta_{\mathrm{core}}^{(i,j)}+\Delta_{\mathrm{spur}}^{(p,q)}.$ With balanced class priors, the corresponding biases yield $m_{i\to j}(\beta^{*})=\tfrac{1}{2}\,\|\Delta_{i,j}\|_{\Sigma},$ so that
$\max_{i}\max_{i\neq j}\,\Phi\!\Big(-\tfrac{1}{2}\,\|\tilde{\Delta}_{i,j}\|_{\Sigma}\Big)\;\leq\;E_{\mathrm{wge}}(\beta^{*})\;\leq\;\max_{i}\sum_{i\neq j}\,\Phi\!\Big(-\tfrac{1}{2}\,\|\tilde{\Delta}_{i,j}\|_{\Sigma}\Big).$
Equivalently, in directional form,
$\|\tilde{\Delta}_{i,j}\|_{\Sigma}=\Big\|\Delta_{\mathrm{core}}^{(i,j)}+\Delta_{\mathrm{spur}}^{(p,q)}\Big\|_{\Sigma}=\Big\|\,\operatorname{cor}(\cdot,\Delta_{\mathrm{core}}^{(i,j)})\,\|\Delta_{\mathrm{core}}^{(i,j)}\|_{\Sigma}+\operatorname{cor}(\cdot,\Delta_{\mathrm{spur}}^{(p,q)})\,\|\Delta_{\mathrm{spur}}^{(p,q)}\|_{\Sigma}\Big\|,$
where $\cdot$” denotes $\beta_{i}^{*}-\beta_{j}^{*}$ normalized in the $\Sigma$-inner product.
We observe that the sign of the spurious term can vary across domain pairs $(p,q)$. A worst-case treatment therefore selects the least favorable sign, yielding
$\min_{(p,q)}\ \operatorname{cor}\!\big(\cdot,\Delta_{\mathrm{spur}}^{(p,q)}\big)\,\big\|\Delta_{\mathrm{spur}}^{(p,q)}\big\|_{\Sigma}\;=\;-\,\max_{(p,q)}\ \big|\operatorname{cor}\!\big(\cdot,\Delta_{\mathrm{spur}}^{(p,q)}\big)\big|\,\big\|\Delta_{\mathrm{spur}}^{(p,q)}\big\|_{\Sigma}.$
Consequently, a conservative lower bound on the pairwise margin is given by
$\underbrace{\operatorname{cor}\!\big(\cdot,\Delta_{\mathrm{core}}^{(i,j)}\big)\,\big\|\Delta_{\mathrm{core}}^{(i,j)}\big\|_{\Sigma}}_{\text{beneficial to increase}}\;-\;\underbrace{\max_{(p,q)}\ \big|\operatorname{cor}\!\big(\cdot,\Delta_{\mathrm{spur}}^{(p,q)}\big)\big|\,\big\|\Delta_{\mathrm{spur}}^{(p,q)}\big\|_{\Sigma}}_{\text{beneficial to decrease}}.$
This decomposition shows that worst-case group-wise margin can be improved by increasing alignment with the core direction while reducing absolute alignment with spurious directions. When combined with the pairwise margin bounds from Theorem 2, this leads to a tighter worst-group error bound by simultaneously raising the lower bound and lowering the upper bound.
Based on this analysis, the SCER framework integrates three key components: (1) Worst-Group Error-based Classification using group-dependent loss weighting, (2) Embedding-based Regularization that identifies and scales core and spurious components through covariance analysis, and (3) Loss Optimization with embedding regularization. This approach effectively mitigates spurious correlations while enhancing class separation, improving robustness under distribution shifts without compromising performance. SCER implements this through regularization terms that control directional correlations, providing a principled approach to robust classification. In conclusion, SCER mitigates spurious correlations by regularizing embedding directions to reduce reliance on spurious features while strengthening core predictive signals. This approach effectively improves worst-group accuracy and enhances robustness under distribution shifts across diverse datasets and modalities.

### v2

@@ Appendix A Methodology
This section provides comprehensive algorithmic details and theoretical proofs for the SCER framework presented in the main paper. We first present detailed step-by-step procedures for SCER implementation, including covariance computation, embedding regularization, and loss optimization, as summarized in the main text. Subsequently, we provide formal proofs for the theoretical foundations underlying our embedding regularization approach, including the decomposition of the embedding space into core and spurious components and the optimality conditions for robustness.
@@ A.1 Overall process of SCER
The Spurious Loss $\mathcal{L}_{\text{spur}}$ penalizes alignment with spurious directions, scaled by the magnitude of spurious variation in the data. The Core Loss $\mathcal{L}_{\text{core}}$ encourages alignment with core directions, with the negative sign ensuring that more substantial core alignment reduces the overall loss. Our loss formulation penalizes alignment with spurious directions and encourages alignment with core directions. By introducing control parameters $\lambda_{\text{spur}}$ and $\lambda_{\text{core}}$, we derive our embedding loss function:
@@
0: Data $(x,y,d)$, parameters $(w,\beta)$, hyperparams $(\eta,\lambda_{\text{spur}},\lambda_{\text{core}})$
@@ A.2 Theoretical Analysis of SCER
We now present a theoretical analysis that formally decomposes the worst-group error. Following prior work (Yao et al., 2022), we adopt a Gaussian mixture model to analyze how worst-group error arises from label-domain interactions.
In the main text, we decompose the classifier $f$ into a feature extractor $f_{w}$ and a linear classifier $f_{\beta}$, where the representation is given by the embedding $x_{\mathrm{emb}}=f_{w}(x)\in\mathbb{R}^{p}$. Throughout this subsection, we conduct the analysis in the embedding space and, with a slight abuse of notation, we overload the symbol $x$ to denote the embedding:
$x\equiv x_{\mathrm{emb}}$
Consequently, all distributions of the form $x\mid(y,d)$ below are to be understood as distributions over embedding vectors.
@@ A.2.1 Binary Setting
Consider a classification problem where each embedding vector $x\in\mathbb{R}^{p}$ is associated with a label $\mathcal{Y}=\{y_{-1},y_{+1}\}$ and a domain $\mathcal{D}=\{d_{R},d_{G}\}$. Each pair $(y,d)\in\mathcal{Y}\times\mathcal{D}$ defines a subpopulation. We assume that the data follows a group-conditional Gaussian distribution $x\mid(y,d)\sim\mathcal{N}(\mu^{(y,d)},\Sigma)$ where $\Sigma$ is positive definite. We further assume that $\Delta_{\text{core}}=\mu^{(y_{+1},d)}-\mu^{(y_{-1},d)}$ is constant across $d$, $\Delta_{\text{spur}}=\mu^{(y,d_{R})}-\mu^{(y,d_{G})}$ is constant across $y$, $\Sigma^{(d_{R})}=\Sigma^{(d_{G})}=\Sigma$, $\mathbb{P}(y=\pm 1)=\mathbb{P}(d\in\{d_{R},d_{G}\})=\frac{1}{2}$, and $\mathbb{P}_{y,d}\approx 0$ for some pairs $(y,d)$ under extreme spurious correlations.
@@
$\log\frac{p(x\mid y_{+1})}{p(x\mid y_{-1})}=(\mu^{(y_{+1})}-\mu^{(y_{-1})})^{\top}\Sigma^{-1}x-\tfrac{1}{2}(\mu^{(y_{+1})}+\mu^{(y_{-1})})^{\top}\Sigma^{-1}(\mu^{(y_{+1})}-\mu^{(y_{-1})}),$
which is linear in $x$ with a normal vector $\Sigma^{-1}(\mu^{(y_{+1})}-\mu^{(y_{-1})})$. Since logistic regression is Fisher-consistent under correct model specification, the cross-entropy minimizer $\beta^{*}$ aligns with the Bayes direction:
$\beta^{*}\propto\Sigma^{-1}(\mu^{(y_{+1})}-\mu^{(y_{-1})}).$
In the extreme shift setting, the effective class means correspond to $\mu^{(y_{+1},d_{R})}$ and $\mu^{(y_{-1},d_{G})}$. Thus
$\beta^{*}\propto\Sigma^{-1}\big(\mu^{(y_{+1},d_{R})}-\mu^{(y_{-1},d_{G})}\big).$
@@
$\mu^{(y_{+1},d_{R})}-\mu^{(y_{-1},d_{G})}=(\mu^{(y_{+1},d_{R})}-\mu^{(y_{-1},d_{R})})+(\mu^{(y_{-1},d_{R})}-\mu^{(y_{-1},d_{G})})=\Delta_{\mathrm{core}}+\Delta_{\mathrm{spur}}=\tilde{\Delta}$,
@@
Remark
Our theoretical analysis relies on modeling the embedding distribution $x\in\mathbb{R}^{p}$ under each subpopulation $(y,d)$ using a Gaussian with a shared covariance matrix $\Sigma$. This assumption enables the tractable LDA-style derivation of Proposition 1 and Theorem 1, where the classifier direction admits the closed-form expression $\beta^{*}\propto\Sigma^{-1}\tilde{\Delta}$. Importantly, this assumption is empirically well-grounded. Seddik et al. (2020) show that deep embeddings exhibit a dominant shared covariance structure across groups, with group-specific differences appearing as higher-order perturbations. Furthermore, the neural collapse phenomenon indicates that within-class covariances converge toward one another as training progresses (Papyan et al., 2020).
@@
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\operatorname{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-{\tfrac{1}{2}\,}\operatorname{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right),$
@@
Using $\beta_{0}=-\beta^{\top}\mathbb{E}[x]$ under extreme spurious correlations, we obtain the following.
Under the extreme shift, the effective training distribution consists of $(y_{+1},d_{R})$ and $(y_{-1},d_{G})$, so $\mathbb{E}[x]=\tfrac{1}{2}(\mu^{(y_{+1},d_{R})}+\mu^{(y_{-1},d_{G})})$ and $\beta_{0}=-\beta^{\top}\mathbb{E}[x]$. Substituting into the subgroup error formula gives $\mathbb{P}(f(x)\neq y\mid(y,d))=\Phi\!\bigl(-\tfrac{y\,\beta^{\top}(\mu^{(y,d)}-\mathbb{E}[x])}{\|\beta\|_{\Sigma}}\bigr)$. For $(y_{+1},d_{R})$: $\mu^{(y_{+1},d_{R})}-\mathbb{E}[x]=\tfrac{1}{2}(\mu^{(y_{+1},d_{R})}-\mu^{(y_{-1},d_{G})})=\tfrac{1}{2}\tilde{\Delta}$, hence $E(y_{+1},d_{R})=\Phi\!\bigl(-\tfrac{\tilde{\Delta}^{\top}\beta}{2\|\beta\|_{\Sigma}}\bigr)$. Applying the same procedure to all four subgroups:
$E(y_{+1},d_{R})$ $=\Phi\!\left(-\tfrac{1}{2}\tfrac{\tilde{\Delta}^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$ (seen)
$E(y_{-1},d_{R})$ $=\Phi\!\left(\tfrac{(\tfrac{1}{2}\tilde{\Delta}-\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$ (unseen)
$E(y_{+1},d_{G})$ $=\Phi\!\left(\tfrac{(\tfrac{1}{2}\tilde{\Delta}-\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}\right),$ (unseen)
$E(y_{-1},d_{G})$ $=\Phi\!\left(-\tfrac{1}{2}\tfrac{\tilde{\Delta}^{\top}\beta}{\|\beta\|_{\Sigma}}\right).$ (seen)
@@
$E(y_{+1},d_{R})=\Phi(-z),\qquad E({y_{-1}},d_{G})=\Phi(-z).$
By Proposition 1, $\beta^{*}\propto\Sigma^{-1}\tilde{\Delta}$, so $\tilde{\Delta}^{\top}\beta^{*}\propto\|\tilde{\Delta}\|_{\Sigma^{-1}}^{2}>0$, implying $z>0$. Both seen subgroups therefore have identical error $\Phi(-z)<\Phi(0)=0.5$. The unseen subgroups $(y_{-1},d_{R}),(y_{+1},d_{G})$ have error $\Phi(w)$ where $w=\tfrac{(\frac{1}{2}\tilde{\Delta}-\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}$. Note that $w=z-\tfrac{\Delta_{\mathrm{core}}^{\top}\beta}{\|\beta\|_{\Sigma}}$, so increasing the core alignment $\Delta_{\mathrm{core}}^{\top}\beta$ reduces $w$ and thus $\Phi(w)$.
$E_{\mathrm{wge}}=\max\{\Phi(-z),\Phi(w)\}.$
Finally, writing $\tilde{\Delta}=\Delta_{\mathrm{spur}}+\Delta_{\mathrm{core}}$ and substituting into $w$ and z-z:
$w=\tfrac{1}{2}\tfrac{(\Delta_{\mathrm{spur}}+\Delta_{\mathrm{core}}-2\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}=+\tfrac{1}{2}\tfrac{\Delta_{\mathrm{spur}}^{\top}\beta}{\|\beta\|_{\Sigma}}-\tfrac{1}{2}\tfrac{\Delta_{\mathrm{core}}^{\top}\beta}{\|\beta\|_{\Sigma}},$
$z=-\tfrac{1}{2}\tfrac{(\Delta_{\mathrm{spur}}+\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}=-\tfrac{1}{2}\tfrac{\Delta_{\mathrm{spur}}^{\top}\beta}{\|\beta\|_{\Sigma}}-\tfrac{1}{2}\tfrac{\Delta_{\mathrm{core}}^{\top}\beta}{\|\beta\|_{\Sigma}}.$
Since $E_{\mathrm{wge}}=\Phi(\max\{-z,w\})$, we obtain
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\frac{\beta^{\top}\Delta_{\mathrm{spur}}}{\|\beta\|_{\Sigma}}-{\tfrac{1}{2}}\frac{\beta^{\top}\Delta_{\mathrm{core}}}{\|\beta\|_{\Sigma}}\right).$
where ++ corresponds to $w$ (unseen group is worst) and - corresponds to z-z (seen group is worst). With Proposition 1 and the identity
@@
The sign $\pm$ appearing in Theorem 1 reflects which subgroup constitutes the worst group. Specifically, since $E_{\mathrm{wge}}=\Phi(\max\{-z,w\})$ where $z=\tfrac{\tilde{\Delta}^{\top}\beta}{2\|\beta\|_{\Sigma}}>0$ and $w=\tfrac{(\frac{1}{2}\Delta_{\mathrm{spur}}-\frac{1}{2}\Delta_{\mathrm{core}})^{\top}\beta}{\|\beta\|_{\Sigma}}$:
  • ++ sign: $\Delta_{\mathrm{spur}}^{\top}\beta^{*}>0$, so w>zw>-z and the unseen subgroups $(y_{-1},d_{R}),(y_{+1},d_{G})$ are the worst group.
  • - sign: $\Delta_{\mathrm{spur}}^{\top}\beta^{*}<0$, so z>w-z>w and the seen subgroups $(y_{+1},d_{R}),(y_{-1},d_{G})$ are the worst group.
By Proposition 1, $\beta^{*}\propto\Sigma^{-1}\tilde{\Delta}$ implies $\Delta_{\mathrm{spur}}^{\top}\beta^{*}\propto\|\Delta_{\mathrm{spur}}\|_{\Sigma^{-1}}^{2}+\Delta_{\mathrm{spur}}^{\top}\Sigma^{-1}\Delta_{\mathrm{core}}>0$ in the typical regime, so the ++ sign (unseen groups worst) applies under ERM with spurious correlations.
Remark
While the extreme correlation assumption is strong, such settings have been widely adopted in prior theoretical studies to isolate and analyze the effect of spurious correlations in a tractable manner (Yao et al., 2022; Lai and Muthukumar, 2024). Following this established tradition, we present our analysis and worst-group error decomposition under these extreme conditions. Empirically, we demonstrate in Section 4 that SCER remains effective even under less extreme correlation regimes, suggesting that the insights derived from this theoretical framework extend to more realistic settings.
@@
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\mathrm{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-{\tfrac{1}{2}\,}\mathrm{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right),$ (4)
@@
Equation 4 shows that two competing components govern the worst-group error:
@@
The first term increases the error by aligning with spurious directions, while the second decreases the error by aligning with core directions. Thus, minimizing $E_{\mathrm{wge}}$ requires simultaneously suppressing the spurious term and enhancing the core term.
@@
In Theorem 1, we derived Equation 4 under the strict ERM setting with Gaussian assumptions, where the optimal $\beta^{*}$ has the closed form $\beta^{*}\propto\Sigma^{-1}\tilde{\Delta}$. In practice, however, $\beta$ corresponds to the weight parameters of a deep neural classifier trained via stochastic optimization. While $\beta$ may not coincide with the ERM minimizer, the decomposition in Equation 4 remains structurally valid: worst-group error is still determined by the trade-off between spurious and core alignment. This observation motivates applying the embedding regularization loss in Equation 5 to deep models.